# Discrete-time thermodynamic speed limit

### v1

@@ III.2 Section speed limit
for any intermediate times $t_{\rm I}$ and $t_{\rm F}$ within $[t_{0},t_{N}]$. Then, from the monotonically increasing property of $h(x)$, the stepwise speed limit (18) can be simplified as
@@
We call Eq. (20) section speed limit. By introducing a function $g(x)\equiv h(x)/(2x)$, one can rearrange Eq. (20) in the following conventional form of a speed limit:
$t_{N}\geq\frac{d_{\textrm{TV}}[\mathbf{P}(t_{\rm F}),\mathbf{P}(t_{\rm I})]}{\bar{A}g^{-1}\left(\frac{\tilde{\Sigma}}{2d_{\textrm{TV}}[\mathbf{P}(t_{\rm F}),\mathbf{P}(t_{\rm I})]}\right)},$ (21)
where $\bar{A}=A_{\textrm{tot}}/t_{N}$. When we set $t_{\rm I}=t_{0}$ and $t_{\rm F}=t_{N}$, Eq. (21) has the same mathematical form as the continuous-time speed limit Delvenne and Falasco (2021); Lee et al. (2021).
It is worth comparing the EP bound derived from our results with the TUR for discrete-time Markov chains. The EP lower bound of the discrete-time TUR is exponentially suppressed Proesmans and Van den Broeck (2017) or proportional to the minimum staying probability Liu et al. (2020). Moreover, the discrete-time TUR applies only to systems without time-dependent driving, whereas our speed limit has no such constraint. This highlights the potential utility of our result for estimating EP in discrete-time processes with arbitrary driving protocols.
In addition, the section speed limit can yield a nontrivial bound for cyclic processes, whereas the conventional speed limit becomes trivial due to the complete cycle condition $\mathbf{P}(t_{0})=\mathbf{P}(t_{N})$ in a cyclic process. We present a concrete application of the section speed limit on a cyclic steady state in Sec. V.
From Eq. (10), the speed limit for the time-backward EP can be obtained using
@@ III.3 Saturation condition
We now investigate the condition for each speed limit being saturated. They saturate when the equality conditions for all inequalities used to derive each speed limit are satisfied. In this regard, inequalities (12) and (III.1) are used to derive the stepwise speed limit, and additional inequalities (17) and (19) are applied for the section speed limit. First, the equality condition of Eq. (III.1) is satisfied when the system has only two states. Interestingly, for a two-state system, Eq. (12) also saturates regardless of the protocol $M_{nm}(t)$ when we adopt $h_{\rm sym}(x)=x\ln[(1+x)/(1-x)]$. Therefore, the stepwise speed limit always saturates for any two-state system. The equal sign of Eq. (17) is attainable when the following condition is satisfied:
@@
for any $i\neq j$. Finding the general saturation condition for inequality (19) is difficult. Nonetheless, the equality of Eq. (19) is attainable when the process ends after only a single time step, i.e., $N=1$, as the section speed limit and stepwise speed limit become identical when $N=1$.
In short, regardless of the protocol $M_{n,m}(t_{i})$, the equality of the stepwise speed limit is achieved for two-state systems with any number of time steps, and the section speed limit saturates in one-time step processes of a system with only two states.
@@ IV Speed limit with unidirectional transitions
We now extend our results to systems that include both bidirectional and unidirectional transitions. A bidirectional transition refers to a transition where the corresponding backward transition exists, while a unidirectional transition refers to a transition where the backward transition is absent. We use a superscript $\alpha={\rm u}$ or ${\rm b}$ to represent the transition probability associated with the bidirectional or unidirectional transition, respectively. That is, $M^{({\rm u})}_{n,m}(t_{i})$ and $M^{({\rm b})}_{n,m}(t_{i})$ denote that the transition from $m$ state to $n$ state is unidirectional and bidirectional, respectively. Therefore, if $M^{({\rm u})}_{n,m}(t_{i})$ is nonzero, $M^{({\rm u})}_{m,n}(t_{i})$ must be 00. Conversely, if $M^{({\rm b})}_{n,m}(t_{i})$ is nonzero, $M^{({\rm b})}_{m,n}(t_{i})$ must also be nonzero. With this notation, the equation of motion can be written as
@@
Since unidirectional transitions violate the microscopic reversibility, they should be excluded from evaluating EP; otherwise, the EP diverges. We denote the EPs for one time step excluding unidirectional transitions as
$\Delta\tilde{\Sigma}^{({\rm b})}(t_{i})\equiv{\sum_{n,m}}^{({\rm b})}M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})}{M_{m,n}^{({\rm b})}(t_{i})P_{n}(t_{i})}$ (25)
and
$\Delta\Sigma^{({\rm b})}(t_{i})\equiv{\sum_{n,m}}^{({\rm b})}M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})}{M_{m,n}^{({\rm b})}(t_{i})P_{n}(t_{i+1})},$ (26)
and we define two kinds of dynamical activities
@@
Additionally, by applying Eqs. (17) and (19) to Eq. (32), we obtain the speed limit with unidirectional transitions as well, which is
@@
We note that there is a speed limit for a continuous-time process with unidirectional transitions in terms of a thermodynamic metric Gupta and Busiello (2020), different from the time-continuum limit of (33).
@@ V Numerical verification
In this section, we numerically verify our results through three examples: a one-time-step process of a two-state system, a cyclic process of a two-state system, and a relaxation process involving a unidirectional transition.
V.1 One-time-step process
Figure 1: Plot of EP versus total dynamical activity for one-time-step relaxation process of the two-state system presented in Sec. V.1. Both the $x$ and $y$ axes are scaled by total variation distance. $+$ and $\times$ symbols represent scaled time-backward EP and time-reversed EP, respectively. The cyan dashed line denotes the speed limit bound estimated from $h(x)=h_{\rm sym}(x)$ function. The parameters used for this plot are as follows: $\mu=T=1$, $0\leq E_{1}\leq 10$, $0\leq\Delta t\leq 20$ and $0.001\leq g\leq 0.99$.
The first example is a one-time-step evolution of a system with two states, labeled as 00 and $1$. The energy gap between these states is denoted as $E_{1}$. To determine the transition probabilities $M_{n,m}$ for a discrete-time process, we begin by examining a continuous-time relaxation process for the two-state system, which is in contact with a thermal reservoir at temperature $T$. The master equation of this continuous-time dynamics is
@@ V.1 One-time-step process
where the transition rates $R_{1,0}=\mu\gamma$ and $R_{0,1}=\mu(1-\gamma)$ follows the Arrhenius rule with $\gamma=e^{-E_{1}/T}/(1+e^{-E_{1}/T})$. The diagonal elements of $\mathbf{R}$ are determined by $R_{0,0}=-R_{1,0}$ and $R_{1,1}=-R_{0,1}$ due to the normalization condition $\sum_{n}R_{n,m}=0$. The formal solution of Eq. (34) is
@@
Figure 1 shows the plot of time-reserved and time-backward EPs with respect to the total dynamical activity. Note that the $x$ and $y$ axes of the plot are scaled by the total variation distance. The parameters $N=T=\mu=1$ are used for this plot. The initial probabilities are parameterized as $P_{0}(t_{0})=g$ and $P_{1}(t_{0})=1-g$. By choosing various values of $g$, $E_{1}$ and $\Delta t$, we evaluate the scaled EPs and $A_{\rm tot}$. The plot demonstrates that time-reversed EP saturates to the speed limit bound estimated from $h(x)=h_{\rm sym}(x)$. This verifies the saturation condition discussed in Sec. III.3. On the other hand, time-backward EP violates the bound.
@@ V.2 Cyclic steady-state process
Figure 2: Plot of time-reversed EP versus total dynamical activity for the cyclic steady-state process presented in Sec. V.2. Both the $x$ and $y$ axes are scaled by total variation distance. The dashed line denotes the section speed limit using $h(x)=h_{\rm sym}(x)$ function. Different symbols represent different choices of $n$ and $m$ for total variation distance $d_{\rm TV}[\mathbf{P}(t_{n}),\mathbf{P}(t_{m})]$. The parameters used for this plot are as follows: $\mu=T=1$, $E_{0}=0$, $0.01\leq\Delta E\leq 7$, and $0.01\leq\Delta\tau\leq 7$.
@@
To verify the section speed limit, we calculate the time-reversed EP and dynamical activity for the complete cycle, while also evaluating the total variation distance for all possible combinations of intermediate distributions at cyclic steady state. For this calculation, we set $\mu=T=1$ and used various values of $\Delta E$ and $\Delta\tau$ within the range from $0.01$ to $7$. The calculation results are presented in Fig. 2. The figure illustrates that for all non-zero total variation distances, the section speed limit provides meaningful bounds on EP. Especially, $d_{\textrm{TV}}[\mathbf{P}(t_{1}),\mathbf{P}(t_{3})]$ gives the tightest bound. This contrasts with the conventional speed limit, determined by the initial and final distributions, which yields a meaningless bound for a cyclic process. Thus, the results demonstrate the usefulness of the derived speed limit in EP estimation problem for a cyclic process.
Figure 3: Plot of time-reversed EP versus dynamical activity of bidirectional transitions for the one-time-step relaxation process of the system, including a unidirectional transition, as presented in Sec. V.3. Both the $x$ and $y$ axes are scaled by $d_{\rm TV}[\mathbf{P}(t_{1}),\mathbf{P}(t_{0})]-A^{({\rm u})}(t_{0})$. The dashed line represents the speed limit for a system including unidirectional transitions, given by Eq. (33), using $h(x)=h_{\rm sym}(x)$. The black dots denote calculation results obtained from various values of $g$, $M_{1,0}$, and $M_{0,1}$ within the range ($0.01,0.9$) under the constraint $d_{\textrm{TV}}[\mathbf{P}(t_{1}),\mathbf{P}(t_{0})]-A^{(u)}(t_{0})>0$.
V.3 System with unidirectional transition

### v2

@@ III.2 Interval speed limit
for any intermediate times $t_{\rm I}$ and $t_{\rm F}$ within $[t_{0},t_{N}]$. Then, from the monotonically increasing property of $h(x)$, a simplified bound can be obtained by
@@
We call Eq. (21) interval speed limit. By using the relation $h(x)=2x\tanh^{-1}x$, one can rearrange Eq. (21) in the following conventional form of a speed limit:
$t_{N}\geq\frac{d_{\textrm{TV}}[\mathbf{P}(t_{\rm F}),\mathbf{P}(t_{\rm I})]}{\bar{A}\tanh\left(\frac{\tilde{\Sigma}}{2d_{\textrm{TV}}[\mathbf{P}(t_{\rm F}),\mathbf{P}(t_{\rm I})]}\right)},$ (22)
where $\bar{A}\equiv A_{\textrm{tot}}/t_{N}$. When we set $t_{\rm I}=t_{0}$ and $t_{\rm F}=t_{N}$, Eq. (22) has the same mathematical form as the continuous-time speed limit Delvenne and Falasco (2024); Lee et al. (2021a). In general cases where the physical time is irrelevant, a bound for the number of steps can be expressed as
$N\geq\frac{d_{\textrm{TV}}[\mathbf{P}(t_{\rm F}),\mathbf{P}(t_{\rm I})]}{\bar{A}_{N}\tanh\left(\frac{\tilde{\Sigma}}{2d_{\textrm{TV}}[\mathbf{P}(t_{\rm F}),\mathbf{P}(t_{\rm I})]}\right)}$ (23)
with the average total activity $\bar{A}_{N}\equiv A_{\rm tot}/N$ per unit transition.
It is worth comparing the EP bound derived from our results with the TUR for discrete-time Markov chains. The EP lower bound of the discrete-time TUR is exponentially suppressed Proesmans and Van den Broeck (2017) or proportional to the minimum staying probability Liu et al. (2020). Moreover, the existing discrete-time TUR applies only to systems without time-dependent driving, whereas our speed limit has no such constraint. This highlights the potential utility of our result for estimating EP in discrete-time processes with arbitrary driving protocols.
In addition, the interval speed limit can yield a nontrivial bound for cyclic processes, whereas the conventional speed limit becomes trivial due to the complete cycle condition $\mathbf{P}(t_{0})=\mathbf{P}(t_{N})$ in a cyclic process. We present a concrete application of the interval speed limit on a cyclic steady state in Sec. V.
Using Eq. (6), one can show that
$\Sigma=\sum_{i=0}^{N-1}\left(\Delta\tilde{\Sigma}(t_{i})-\Delta\Sigma^{\prime}(t_{i})\right)=\tilde{\Sigma}-\sum_{i=0}^{N-1}\Delta\Sigma^{\prime}(t_{i}).$ (24)
It yields the speed limit for the time-backward EP given by
@@ III.3 Saturation condition
We now investigate the condition for each speed limit being saturated. They saturate when the equality conditions for all inequalities used to derive each speed limit are satisfied. In this regard, inequalities (12) and (III.1) are used to derive the stepwise speed limit, and additional inequalities (18) and (20) are applied for the interval speed limit . First, the equality condition of Eq. (III.1) is satisfied when the system has only two states. Interestingly, for a two-state system, Eq. (12) also saturates regardless of the protocol $M_{n,m}(t)$. Therefore, the stepwise speed limit always saturates for any two-state system. The equal sign of Eq. (18) is attainable when the following condition is satisfied:
@@
for any $i\neq j$. Finding the general saturation condition for inequality (20) is difficult. Nonetheless, the equality of Eq. (20) is trivially satisfied when the process ends after only a single time step, i.e., $N=1$, as the interval speed limit and stepwise speed limit become identical when $N=1$.
In short, regardless of the protocol $M_{n,m}(t_{i})$, the equality of the stepwise speed limit can be achieved for two-state systems with any number of time steps, and the interval speed limit saturates in one-time step processes of a system with only two states.
@@ IV Speed limit with unidirectional transitions
We now extend our results to systems that involve both bidirectional and unidirectional transitions. A bidirectional transition refers to a transition where the corresponding backward transition exists, while a unidirectional transition refers to a transition where the backward transition is absent. We use a superscript $\alpha={\rm u}$ or ${\rm b}$ to represent the transition probability associated with the unidirectional or bidirectional transition, respectively. That is, $M^{({\rm u})}_{n,m}(t_{i})$ and $M^{({\rm b})}_{n,m}(t_{i})$ denote that the transition from $m$ state to $n$ state is unidirectional and bidirectional, respectively. Therefore, if $M^{({\rm u})}_{n,m}(t_{i})$ is nonzero, $M^{({\rm u})}_{m,n}(t_{i})$ must be 00. Conversely, if $M^{({\rm b})}_{n,m}(t_{i})$ is nonzero, $M^{({\rm b})}_{m,n}(t_{i})$ must also be nonzero. With this notation, the equation of motion can be written as
@@
Defining EP for a system with unidirectional transitions is a nontrivial problem. Thus, several approaches have been suggested to extend the results of EP for Markovian systems with only bidirectional transitions to systems with unidirectional transitions Rahav and Harbola (2014); Fuchs et al. (2016); Gupta and Busiello (2020); Busiello et al. (2020); Mori et al. (2023); Manzano et al. (2024). This difficulty arises because directly applying the EP definition in Eqs. (4) or (5) to systems with unidirectional transitions leads to divergences. A similar problem also arises in systems with absolutely irreversible processes Murashita et al. (2014); Hoang et al. (2016). One possible resolution is to define EP solely based on bidirectional transitions. This definition has been adopted in studies of fluctuation theorems for systems with unidirectional transitions Rahav and Harbola (2014), as well as in other studies on continuous-time Markov processes with unidirectional transitions Fuchs et al. (2016); Busiello et al. (2020); Pal et al. (2021a); Manzano et al. (2024).
This definition of entropy production can be obtained by decomposing the system entropy change as
$\Delta S_{\rm sys}(t_{i})=\Delta S^{\rm(b)}_{\rm sys}(t_{i})+\Delta S^{\rm(u)}_{\rm sys}(t_{i}),$ (28)
where
$\Delta S^{(\alpha)}_{\rm sys}(t_{i})\equiv{\sum_{n,m}}^{(\alpha)}M_{n,m}^{(\alpha)}(t_{i})P_{m}(t_{i})\ln{\frac{P_{m}(t_{i})}{P_{n}(t_{i+1})}}$ (29)
represents the entropy change associated with transition type $\alpha$. The first term, corresponding to the bidirectional transitions, can be further decomposed as Fuchs et al. (2016); Gupta and Busiello (2020); Busiello et al. (2020):
$\Delta S^{\rm(b)}_{\rm sys}(t_{i})=$ $\Delta\Sigma^{({\rm b})}(t_{i})-{\sum_{n,m}}^{({\rm b})}M_{n,m}^{\rm(b)}(t_{i})P_{m}(t_{i})\ln{\frac{M^{({\rm b})}_{n,m}(t_{i})}{M^{({\rm b})}_{m,n}(t_{i})}},$ (30)
where the time-backward EP for bidirectional transitions is defined as
$\Delta\Sigma^{({\rm b})}(t_{i})\equiv{\sum_{n,m}}^{({\rm b})}M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})}{M_{m,n}^{({\rm b})}(t_{i})P_{n}(t_{i+1})}.$ (31)
This decomposition shows that the modified EP, $\Delta\Sigma^{\rm(b)}$, can be interpreted as the entropy production arising solely from the bidirectional transitions. Similarly, the time-reversed EP only for bidirectional transitions can be defined as
$\Delta\tilde{\Sigma}^{({\rm b})}(t_{i})\equiv$ ${\sum_{n,m}}^{({\rm b})}M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}^{({\rm b})}(t_{i})P_{m}(t_{i})}{M_{m,n}^{({\rm b})}(t_{i})P_{n}(t_{i})},$ (32)
and, in terms of the time-reversed EP, the system entropy change is decomposed as
$\Delta S_{\rm sys}(t_{i})=$ $\Delta\tilde{\Sigma}^{({\rm b})}(t_{i})+{\sum_{n,m}}^{({\rm u})}M_{n,m}^{\rm(u)}(t_{i})P_{m}(t_{i})\ln{\frac{P_{m}(t_{i})}{P_{n}(t_{i+1})}}$
$-{\sum_{n,m}}^{({\rm b})}M_{n,m}^{\rm(b)}(t_{i})P_{m}(t_{i})\ln{\frac{M^{({\rm b})}_{n,m}(t_{i})P_{m}(t_{i})}{M^{({\rm b})}_{m,n}(t_{i})P_{n}(t_{i})}}.$ (33)
The physical interpretation of EPs defined in Eqs. (31) and (32) depends on the specific context in which the unidirectional transitions are introduced. For instance, in systems with stochastic resetting, this EP corresponds to the EP generated during diffusion between successive resetting events Fuchs et al. (2016). In general, such EP quantifies the degree of irreversibility arising solely from bidirectional transitions. In the following, we demonstrate that using this definition allows us to derive a speed limit.
Two types of dynamical activities can be defined as
@@
Figure 1: Plot of EP versus total dynamical activity for one-time-step relaxation process of the two-state system presented in Sec. V.1. Both the $x$ and $y$ axes are scaled by the total variation distance between the initial and final state probabilities for each point. $+$ and $\times$ symbols represent scaled time-backward EP and time-reversed EP, respectively. The cyan dashed line denotes the speed limit bound estimated from $h(x)$ function. The parameters used for this plot are as follows: $\mu=T=1$, $0\leq E_{1}\leq 10$, $0\leq\Delta t\leq 20$ and $0.001\leq g\leq 0.99$.
Additionally, by applying Eqs. (18) and (20) to Eq. (39), we obtain the interval speed limit with unidirectional transitions as well, which is
@@
We note that there is a speed limit for a continuous-time process with unidirectional transitions in terms of a thermodynamic metric Gupta and Busiello (2020), which differs from the time-continuum limit of (40).
@@ V Numerical verification
In this section, we numerically verify our results through three toy models: the relaxation process of a two-state system, a cyclic process of a two-state system, and a relaxation process involving a unidirectional transition. Additionally, we apply our results to cell-state dynamics of human breast cancer and deterministic finite automaton for identifying the multiples of 4.
V.1 Relaxation process of a two-level system
Figure 2: Plot of the right-hand side of Eq. (23) versus the number of steps, denoted as $N$, for the $N$-step relaxation process of the two-state system presented in Sec. V.1. $+$ and $\times$ symbols represent the bound on the number of steps in terms of the time-backward EP without the corrections term, $\sum_{i=0}^{N-1}\Delta\Sigma^{\prime}(t_{i})$ and the time-reversed EP, respectively. The cyan dashed line denotes $y=x$ function. The parameters used for this plot are the same as those in the previous figure.
The first example considers a relaxation process with two states, labeled as 00 and $1$. The energy gap between these states is denoted as $E_{1}$. To determine the transition probabilities $M_{n,m}$ for a discrete-time process, we begin by examining a continuous-time relaxation process for the two-state system, which is in contact with a thermal reservoir at temperature $T$. The master equation of this continuous-time dynamics is
@@ V.1 Relaxation process of a two-level system
where the transition rates $R_{1,0}=\mu\gamma$ and $R_{0,1}=\mu(1-\gamma)$ follows the Arrhenius rule with $\gamma=e^{-E_{1}/T}/(1+e^{-E_{1}/T})$. The diagonal elements of $\mathbf{R}$ are determined by $R_{0,0}=-R_{1,0}$ and $R_{1,1}=-R_{0,1}$ due to the normalization condition $\sum_{n}R_{n,m}=0$. The system’s probability vector is defined as $\mathbf{P}(t)=[P_{0}(t),P_{1}(t)]^{\textsf{T}}$. The formal solution of Eq. (41) is
@@
Figure 1 shows the plot of time-reversed and time-backward EPs with respect to the total dynamical activity. The data were generated from the one-step relaxation process of the two-state system. Note that the $x$ and $y$ axes of the plot are scaled by the total variation distance for each point. The parameters $N=T=\mu=1$ are used for this plot. The initial probabilities are parameterized as $P_{0}(t_{0})=g$ and $P_{1}(t_{0})=1-g$. By choosing various values of $g$, $E_{1}$ and $\Delta t$, we evaluate the scaled EPs and $A_{\rm tot}$. The plot demonstrates that time-reversed EP saturates to the speed limit bound. This verifies the saturation condition discussed in Sec. III.3. On the other hand, the results for time-backward EP, represented by red cross symbols, are located below the bound, indicating that time-backward EP violates Eq. (21).
To verify Eq. (23), we also calculate the lower bounds in $N$-step relaxation processes for $1\leq N\leq 7$. Figure 2 shows the results, illustrating that for $N=1$, the time-reversed EP bound exactly matches $N$, in agreement with theoretical predictions. For $N\leq 2$, the time-backward EP bound, obtained by replacing $\tilde{\Sigma}$ with $\Sigma$, occasionally exceeds $N$, indicating a violation of the conventional speed limit. However, such violations become rare for $N>2$. This observation arises because, in most processes, the time-reversed EP bound lies quite below $N$. Only in slowly evolving processes, the time-reversed EP bound closely approaches $N$, but in this case $\Delta\Sigma^{\prime}$ is too small to induce a violation.
@@ V.2 Cyclic steady-state process
Figure 3: Plot of time-reversed EP versus total dynamical activity for the cyclic steady-state process presented in Sec. V.2. Both the $x$ and $y$ axes are scaled by the total variation distance between the initial and final state probabilities for each point. The dashed line denotes the interval speed limit. Different symbols represent different choices of $n$ and $m$ for total variation distance $d_{\rm TV}[\mathbf{P}(t_{n}),\mathbf{P}(t_{m})]$. The parameters used for this plot are as follows: $\mu=T=1$, $E_{0}=0$, $0.01\leq\Delta E\leq 7$, and $0.01\leq\Delta\tau\leq 7$.
@@
To verify the interval speed limit, we calculate the time-reversed EP and dynamical activity for the complete cycle, while also evaluating the total variation distance for all possible combinations of intermediate distributions at cyclic steady state. For this calculation, we set $\mu=T=1$ and used various values of $\Delta E$ and $\Delta\tau$ within the range from $0.01$ to $7$. The calculation results are presented in Fig. 3. The figure illustrates that for all non-zero total variation distances, the interval speed limit provides meaningful bounds on EP. Especially, $d_{\textrm{TV}}[\mathbf{P}(t_{3}),\mathbf{P}(t_{1})]$ yields the result closest to the bound. This contrasts with the conventional speed limit, determined by the initial and final distributions, which yields a meaningless bound for a cyclic process. Thus, the results demonstrate the usefulness of the interval speed limit in EP estimation problem for a cyclic process.
Figure 4: Plot of time-reversed EP versus dynamical activity of bidirectional transitions for the one-time-step relaxation process of the system, including a unidirectional transition, as presented in Sec. V.3. Both the $x$ and $y$ axes are scaled by the total variation distance minus unidirectional activity, $d_{\rm TV}[\mathbf{P}(t_{1}),\mathbf{P}(t_{0})]-A^{({\rm u})}(t_{0})$, for each point. The dashed line represents the speed limit for a system including unidirectional transitions, given by Eq. (40). The black dots denote calculation results obtained from various values of $g$, $M_{1,0}$, and $M_{0,1}$ within the range ($0.01,0.9$) under the constraint $d_{\textrm{TV}}[\mathbf{P}(t_{1}),\mathbf{P}(t_{0})]-A^{(u)}(t_{0})>0$.
V.3 Three-level system with unidirectional transition